For the rest of the problems involving this pendul
Answer #1

Here ,

m = 20 Kg

l = 20 m

initial speed at the top , u = 12 m/s

at the bottom of the path

let the speed of pendulum is v

Using conseravtion of energy

0.5 * m * v^2 = 0.5 * m * u^2 + m *g * (2 * l)

0.5 * v^2 = 0.5 * 12^2 + 2 * 9.8 * 20

solving for v

v = 30.5 m/s = 31 m/s

the speed of the ball at the bottom is 31 m/s

20)

for u = 16 m/s

Using conseravtion of energy

0.5 * m * v^2 = 0.5 * m * u^2 + m *g * (2 * l)

0.5 * v^2 = 0.5 * 16^2 + 2 * 9.8 * 20

solving for v

v = 32 m/s = 32 m/s

the speed of the ball at the bottom is 32 m/s